Consider the following financial data.

US 10-Year T-Bond Yield = 1.5%

Market Risk Premium = 6.25%

Tax Rate = 21%

Also the following data for Ford Motor Company:

Stock Price = $9.17

Market Cap = $35.946 billion

Beta = .95

Moodys = A1 (115 basis points)

Total Debt = $156.06 billion

Number of Shares Outstanding = 3.92 billion

EPS = $.54

Return on Assets = .93%

Total Debt/Equity (Book Value) = 431.88

Book Value/share = $9.05

Revenues = $158.65 billion

Price/Earning Multiple = 16.98

Calculate the Cost of Capital AND MVA for Ford Motor Company.
Answer #1

Cost of capital as per CAPM equation = Risk free rate + Beta * Market risk premium

= 1.50% + 0.95 * 6.25% = 7.438%

Market value added = Market value of equity + market value of debt - book value of equity - book value of debt

Book value of equity = Shares outstanding * Book value per share

= 3.92 * 9.05 = 35.476 Billion

Book value of debt = Book value of equity * Debt / equity ratio

= 35.476 * 431.88 /100 = 153.214 Billion

Market value added = 35.946 + 156.06 - 35.476 - 153.214

= 3.316 Billion